OpenStudy (anonymous):

let f(x)=(x^2)(e^5x) find the derivative f'(x) and the second derivative, f"(x)

OpenStudy (anonymous):

DERIVATE OF 1ST TIMES 2ND PLUS DERIVATIVE OF 2ND TIMES 1ST

OpenStudy (anonymous):

PRODUCT RULE

OpenStudy (anonymous):

Use the product rule! f(x) = (x^2)(e^5x) First times the derivative of the second, plus, second times the derivative of the first: (d/dx) f(x) = (x^2) d/dx(e^5x) + (e^5x) d/dx (x^2) what's the derivative of e^5x? well you have to chain rule for that: d/dx (e^5x) = (e^5x) (5) now what is the derivative of x^2? yup! just 2 d/dx (x^2) = 2 so what's f'(x)? Let's just plug in our derivatives f'(x) = (x^2) (e^5x) (5) + (e^5x)(2) tada!

OpenStudy (anonymous):

f'(x) = 2x * e^(5x) + ( 1/5) x² e^(5x) f"(x) = [ 2e^(5x) + (2/5)x e^(5x) ]+ [(2/5)x e^(5x) + [(1/25) x² e^(5x) ]

OpenStudy (anonymous):

-> f"(x) = e^(5x) [ (x²/ 25) + (4x/5) + 2 ]